# Reaction-diffusion models of invasive tree pest spread: quantifying the expansion of oak processionary moth in the UK

### v1

@@ Abstract.
UK woodlands, forests, and urban treescapes are under threat from invasive species, exacerbated by climate change, trade, and transport. Invasive tree pests debilitate their host and disrupt forest ecosystems, thus it is imperative to quantitatively model and predict their spread. Addressing this, we represent the spatial distribution of the pest as a population density field which evolves according to a spatiotemporal reaction-diffusion equation. We solve this intractable system of equations numerically and, from the solution, we determine first arrival times of the pest at locations in the field. The adopted model permits us to obtain the expansion rate of pest spread directly from the model parameters, which we infer in the Bayesian paradigm, using a Markov chain Monte Carlo scheme. We apply our framework to the ongoing spread of oak processionary moth in the UK, an outbreak which continues to grow despite management efforts. We demonstrate that our approach effectively captures the spread of the pest and that this has occurred at a non-constant expansion rate. The proposed framework is a powerful tool for quantitatively modelling the spread of an invasive tree pest and could underpin future prediction and management approaches.
@@ 1        I ntroduction
Invasive tree pests pose a significant ecological, economical, and epidemiological threat both in the UK and globally [1, 2]. The UK hosts at least 121 pests of native tree species which were either introduced or have uncertain origin, and management of the most expensive six is estimated to cost around £919.9 million per year [1]. Climate change is anticipated to favour successful establishment of satellite populations and international trade is known to be a key driver of these introductions [3]. In particular, imports of live oak led to the successful establishment of the oak processionary moth (OPM) in the South of England in 2006 [4]. OPM caterpillars – which defoliate and weaken trees – develop urticating setae which can shed and become airborne [5]. Human and/or animal contact with the caterpillars, their detached fibrous hairs, or their nests’ detritus can lead to cases of caterpillar dermatitis [5, 6, 7]. Inhalation of the setae, which can break apart into microscopic fibres, can lead to respiratory irritation [7].
@@ 2        M aterials and Methods
Here we describe the adopted approach of this paper. In Section 2.1, we present the OPM observational data used to train our reaction-diffusion models; these observed sites are shown in Figure 1. In Section 2.2, we describe the reaction-diffusion model of OPM spread – the FKPP equation – along with our assumptions on the population. In Section 2.3, we define the two models of OPM spread proposed in this work. In Section 2.4, we outline the adopted Bayesian approach to inferring the parameters of these proposed models. Finally, in Section 2.5, we provide the metric used to validate these models.
Figure 1. The distribution of OPM nest observations (see Section 2.1 for details). Each of the $414$ coloured squares indicates a $1\text{km}^{2}$ area where OPM is observed; the colour indicates the earliest year OPM was observed in that area. Black crosses indicate the centre of the initial distribution. Panel (a) shows the distribution on a map of the UK. Panel (b) shows the $256\text{km}^{2}$ grid $\Omega$ adopted as the modelling domain (see Section 2.1), indicated by a square outline on panel (a). Panel (c) shows the approximation of the initial (observed in 2006) distribution. The minor ticks indicate the 11km resolution grid. For the background maps of the UK, we use the shapefile for Q. robur (English oak) provided by [26], normalised to one.
@@ 2.1     O PM observational data
These data are derived from amalgamated survey data (provided by the Forestry Commission and the University of Southampton), obtained from yearly surveys conducted between $2006$ and $2023$ [27]. The adopted reaction-diffusion model (see Section 2.2) captures the first arrival of OPM, but cannot capture the complex spatiotemporal processes behind the wavefront. Thus, we thin the dataset to capture this observed first arrival using a convex layer decomposition. Beginning with the final survey year $t=2023$, we calculate the convex hull of (the locations of) the corresponding observed sites and interpret the vertex set of the hull as the observed wavefront in year $t$. We compare the sites observed in previous years to the observed wavefront and remove any that are located ahead of the wavefront. We set $t\leftarrow t-1$ and repeat until (and including) the initial year $t=2006$. Then, beginning at $t=2006$, we construct the observed wavefront and compare this wavefront to sites observed in later years, removing those located behind the wavefront. We set $t\leftarrow t+1$ and repeat until (and including) $t=2023$. This produces ‘layered’ data that captures the spatial trend of population spread over time – see Figure 2 for the resulting observed distributions for $2014$ and $2023$.
Figure 2. The distributions of observed sites for 2014 (panel (a)) and 2023 (panel (b)), obtained by applying a convex layer decomposition (Section 2.1). For the background maps of the UK, we use the shapefile for Q. robur (English oak) provided by [26], normalised to one.
@@ 2.2     F KPP equation for tree pest spread
with growth rate $r$, diffusivity $D$, and carrying capacity $K$ [28]. The first term in Equation (2.1) describes diffusion of $N$ in space; diffusion represents the dispersal of the population. The second term describes the logistic growth of $N$ in time, where $r$ is the growth rate and $K$ the carrying capacity; logistic growth represents reproduction of the population subject to resource constraints encoded by $K$, the limit of the (local) population density. For constants $r$, $D$, and $K$ (the so-called isotropic FKPP equation), one-dimensional solutions evolve to a propagating wavefront which travels at a speed of
@@
Equation (2.1) has a natural timescale of $1/r$ and lengthscale $\ell$ satisfying $D=\ell^{2}r$ (see e.g. [30]). The natural lengthscale is the OPM dispersal distance, so we set $\ell=0.5$km, given in [11] as the expected distance between a new nest and its parent nest. The wavespeed $v$ in Equation (2.2) can be interpreted as the expansion rate of the population and has units of km/year. Since $\ell=0.5$, we have $v=r$ km/year (with $r>0$, assumed implicitly in what follows). The specific forms we adopt for the growth rate $r$ are given in Section 2.3.
@@ 2.3     P roposed expansion rate(s)
for some real value $r>0$. The parametrisation of $M_{1}$ is $\theta_{1}=(r)$, where we adopt vector notation for consistency (see definition of $\theta_{2}$ below). This monophasic model $M_{1}$ encodes the assumption that the spread of OPM occurs at a constant-rate $r$ throughout the period $2006$ to $2023$. The second model $M_{2}$ is of the form Equation (2.1) with
@@
where $r_{1},r_{2}>0$ are real values and the transition time $t_{c}$ satisfies $2006\leq t_{c}\leq 2023$. Hence, $M_{2}$ is parametrised by $\theta_{2}=(r_{1},r_{2},t_{c})$. This biphasic model $M_{2}$ encodes the assumption that the spread of OPM occurs at a non-constant rate; at a rate of $r_{1}$ from $2006$ up to and including $t_{c}$ and at a rate $r_{2}$ from $t_{c}$ to $2023$. We provide snapshots of a prototypical numerical solution to this model in Figure 3.
@@ 2.4     P arameter inference
The MCMC scheme (Gibbs sampler) described above is iterated until $m$ samples are obtained. We discard the initial $m_{B}\geq p$ samples as burn-in, and informally assess convergence of the remaining $m=m_{T}-m_{B}$ samples via visual diagnostics such as trace plots [25]. Additional scaling of the proposal variance to optimise acceptance rates can be carried out post hoc, if desired [38]. See Appendix A for more details.
@@ 2.5     M odel validation
We run the MCMC scheme for each of the models $M_{i}$ described above. To compare the models, we use the deviance information criterion (DIC), which measures relative quality of models [39]. For a model $M_{i}$ (parametrised by $\theta_{i}$), the DIC is given by
@@
where $D(\theta_{i})=-2\log(\pi(\mathcal{D}|\theta_{i}))$ is the deviance of $\theta_{i}$; this criterion quantifies goodness of fit while penalising over-fitting. We can readily compute $\bar{\theta}_{i}$ and $\overline{D(\theta_{i})}$ using the samples of $\theta$, obtained from the (final) MCMC run. Models with a smaller DIC should be preferred to models with a larger DIC. Finally, we assess goodness of fit of the model with lowest DIC using posterior predictive sampling. Let $\tilde{t}_{k}$ denote a first arrival time at site $\mathbf{s}_{k}$. The predictive density is given by
@@ 3        R esults
Recall that Model 1 is parametrised by $\theta_{1}=(r)$ and $\sigma_{1}$. We initialise the MCMC scheme with values $\theta_{1,0}=4.0$ and $\sigma_{1,0}=2.0$, chosen arbitrarily. After burn-in and thinning the sample output to reduce auto-correlation, we obtain $m_{1}=10^{4}$ samples from the posterior distribution of the parameters of Model 1. This output is summarised by Figures 4 and 5, with the former suggesting satisfactory convergence of the MCMC scheme. In Table 2, we provide a summary of the inferred parameters for each model. For the constant-rate Model 1, we obtain a posterior mode of $\hat{r}=2.91$ with $95\%$ (equi-tailed) credible interval (CI) $(2.77,3.06)$ for the growth parameter $r$. For the observed noise parameter $\sigma$, the posterior mode is $\hat{\sigma}_{1}=4.85$ with $95\%$ CI $(4.54,5.21)$. The DIC reported for this model is $\text{DIC}_{1}=-1728.07$ (2dp).
@@ 4.1     T he spread of oak processionary moth in the UK
The values for the DICs of Models 1 and 2 (Table 2) indicate that Model 2 is the more appropriate choice for these observed data. This suggests there is evidence that the expansion of oak processionary moth in the UK from 2006 to 2023 occurred at a non-constant rate. This conclusion aligns with previous results in the literature [8].
@@
Our evidence of biphasic expansion aligns with previous results, which also indicated two phases of expansion (at rates $1.66$ and $6.17$, respectively) with transition time $2014$ [8]. We note that the weighted average of $\hat{r}_{1}$ and $\hat{r}_{2}$ ($3.28$) is reasonably close in value to the Model 1 growth rate ($2.91$), but is closer to the mean of the previous biphasic estimates ($3.27$) [8]. Since the latter estimate covers $2006$ to $2019$, this could suggest that the expansion rate of OPM has continued to increase from $2019$ to $2023$. This is further supported by the increase in the phase 2 growth rate ($\hat{r}_{2}=6.21$ vs $6.17$). Nevertheless, we note that $6.17$ is within the 95CI of $r_{2}$, so it is possible the rate has stayed the same or decreased. In addition, methodological differences could explain these discrepancies.
@@ 4.2     M odelling approach
Previous estimates of the expansion rate of OPM in the UK were most recently calculated using a maximum distance method [8]. With our reaction-diffusion approach, we can obtain an estimate of the rate directly from our model parameters. This offers benefits; this framework can readily estimate future expansion rates. However, the equation for the expansion rate (Equation (2.2)) is obtained analytically, hence it will be susceptible to error introduced by the choice of numerical integration method. Fortunately, testing indicates acceptable convergence of the numerical scheme for our model, spatial-, and temporal-scales.
Our spatially-explicit modelling approach provides an estimate of the arrival of OPM at every $1\text{km}$ area in a $256\text{km}^{2}$ spatial domain and can be readily extended to larger domain sizes, provided adequate computational resources are available. Our reaction-diffusion equation describes the evolution of the population density field; we apply a binary mask to this to facilitate comparison of our model with observed presence/absence data. Additionally, this could facilitate direct comparison of our approach with similar (compartmental) approaches [18, 19]. Furthermore, the population density field evolves with a clear outward-travelling wavefront (as in Figure 3). This captures the spatial trend of first arrival of the pest without incurring the computational cost of simulating a large number of individuals, for example. The use of a binary mask also allows our model to somewhat account for potential time lag in observation. Time lag is posited to reflect the time needed for populations to reach detectable densities [43].
@@ 4.3     F urther Work
This work was completed as part of JPM’s PhD, supported by the Natural Environment Research Council (NERC) (Grant Number [NE/S007512/1]). LEW acknowledges support from a NERC Knowledge Exchange Fellowship (Grant Number [NE/X000478/1]). We thank Julia Branson from the University of Southampton and Andrew Hoppit from Forestry Commission England for useful discussions on the application to OPM. We also thank Kaitlyn N. Ries for helpful feedback on the manuscript.
@@ Appendix A Inferring model parameters
In step 3, the full conditional for $\theta$ is intractable, hence we adopt a random walk Metropolis-Hastings step to sample from the corresponding distribution. Define $\lambda=\log(\theta)$, where the logarithm is applied component-wise. In step 3 of the Gibbs sampler, we propose a new value $\lambda^{*}=\lambda+{\omega}$, where ${\omega}\sim N({0},\Sigma)$, for a strictly positive-definite innovation matrix $\Sigma$ - the covariance matrix of the proposal distribution. The choice of innovation matrix $\Sigma$ influences the mixing of the Markov chain. Practitioners often adopt an adaptive MCMC scheme wherein the innovation matrix $\Sigma$ is updated at fixed intervals to improve mixing [36]. We adopt the adaptive approach of Haario-et-al [36]. The MCMC scheme is instantiated with an arbitrary, strictly positive-definite, diagonal initial innovation matrix $\Sigma=\Sigma^{0}$ (we use the diagonal of the prior covariance matrix with zero covariances). We select a tuning period $p\ll m$ and a tuning interval $l\ll p$ where, for numerical convenience, we assume $p$ is divisible by $l$. Let $\theta^{(j)}$ denote the $j$th value in the chain. For the first $p$ samples, we update the innovation matrix at every $l^{\text{th}}$ sample. Hence the innovation matrix for the $j^{\text{th}}$ step of the MCMC scheme is given by
@@
where the scaling parameter $s_{d_{k}}=(2.38^{2})/{d_{k}}$ is a heuristic taken from [37] and $d_{k}$ is the dimension of the parameter space of model $M_{k}$. For a justification of this adaptive scheme, we refer the reader to [36].

### v2

@@ Abstract.
UK woodlands, forests, and urban treescapes are under threat from invasive species, exacerbated by climate change, trade, and transport. Invasive tree pests debilitate their host and disrupt forest ecosystems, thus it is imperative to quantitatively model and predict their spread. Addressing this, we model the spread of an invasive pest using a spatiotemporal reaction-diffusion equation, representing the spatial distribution as a population density field. We solve this intractable equation numerically and, from the solution, we determine first arrival times of the pest at locations in the field. The adopted model permits us to obtain the expansion rate of pest spread directly from the model parameters, which we infer in the Bayesian paradigm, using a Markov chain Monte Carlo scheme. We apply our framework to the ongoing spread of oak processionary moth in the UK, an outbreak which continues to grow despite management efforts. We demonstrate that our approach effectively captures the spread of the pest and that this has occurred at a non-constant expansion rate. The proposed framework is a powerful tool for quantitatively modelling the spread of an invasive tree pest and could underpin future prediction and management approaches.
@@ 1        I ntroduction
Invasive tree pests pose a significant ecological, economical, and epidemiological threat both in the UK and globally [1, 2]. The UK hosts at least 121 pests of native tree species which were either introduced or have uncertain origin [3], and management of the most expensive six is estimated to cost around £919.9 million per year [1]. Climate change is anticipated to favour successful establishment of satellite populations and international trade is known to be a key driver of these introductions [3, 4]. In particular, imports of live oak led to the successful establishment of the oak processionary moth (OPM) in the South of England in 2006 [5]. OPM caterpillars – which defoliate and weaken trees – develop urticating setae which can shed and become airborne [6]. Human and/or animal contact with the caterpillars, their detached fibrous hairs, or their nests’ detritus can lead to cases of caterpillar dermatitis [6, 7, 8]. Inhalation of the setae, which can break apart into microscopic fibres, can lead to respiratory irritation [8].
@@ 2        M aterials and Methods
Here we describe the adopted approach of this paper. In Section 2.1, we present the OPM observational data used to train our reaction-diffusion models; these observed sites are shown in Figure 1. In Section 2.2, we describe the reaction-diffusion model of OPM spread – the FKPP equation – along with our assumptions on the population. In Section 2.3, we define the two models of OPM spread proposed in this work. In Section 2.4, we outline the adopted Bayesian approach to inferring the parameters of these proposed models. Finally, in Section 2.5, we provide the metrics used to validate these models.
Figure 1. The distribution of OPM nest observations (see Section 2.1 for details). Each of the $414$ coloured squares indicates a $1\text{km}^{2}$ area where OPM is observed; the colour indicates the earliest year OPM was observed in that area. Black crosses indicate the centre of the initial distribution. Panel (a) shows the distribution on a map of the UK. Panel (b) shows the $256\text{km}^{2}$ grid $\Omega$ adopted as the modelling domain (see Section 2.1), indicated by a square outline on panel (a). Panel (c) shows the approximation of the initial (observed in 2006) distribution. The minor ticks indicate the 11km resolution grid. For the background maps of the UK, we use the shapefile for Q. robur (English oak) provided by [27] (available at [28]), normalised to one.
@@ 2.1     O PM observational data
These data are derived from amalgamated survey data (provided by the Forestry Commission and the University of Southampton), obtained from yearly surveys conducted between $2006$ and $2023$ [29]. The adopted reaction-diffusion model (see Section 2.2) captures the first arrival of OPM, but cannot capture the complex spatiotemporal processes behind the wavefront. Thus, we thin the dataset to capture this observed first arrival using a convex layer decomposition [30]. Beginning with the final survey year $t=2023$, we calculate the convex hull of (the locations of) the corresponding observed sites and interpret the vertex set of the hull as the observed wavefront in year $t$. We compare the sites observed in previous years to the observed wavefront and remove any that are located ahead of the wavefront. We set $t\leftarrow t-1$ and repeat until (and including) the initial year $t=2006$. Then, beginning at $t=2006$, we construct the observed wavefront and compare this wavefront to sites observed in later years, removing those located behind the wavefront. We set $t\leftarrow t+1$ and repeat until (and including) $t=2023$. This produces ‘layered’ data that captures the spatial trend of population spread over time – see Figure 2 for the resulting observed distributions for $2014$ and $2023$.
Figure 2. The distributions of observed sites for 2014 (panel (a)) and 2023 (panel (b)), obtained by applying a convex layer decomposition (Section 2.1). For the background maps of the UK, we use the shapefile for Q. robur (English oak) provided by [27], (available at [28]) normalised to one.
@@ 2.2     F KPP equation for tree pest spread
with growth rate $r$, diffusivity $D$, and carrying capacity $K$ [31]. The first term in Equation (2.1) describes diffusion of $N$ in space; diffusion represents dispersal of the population. The second term describes the logistic growth of $N$ in time, where $r$ is the growth rate and $K$ the carrying capacity; logistic growth represents reproduction of the population subject to resource constraints encoded by $K$, the limit of the (local) population density. For constants $r$, $D$, and $K$ (the so-called isotropic FKPP equation), one-dimensional solutions evolve to a propagating wavefront which travels at a speed of
@@
Equation (2.1) has a natural lengthscale $\ell$ satisfying $D=\ell^{2}r$ (see e.g. [33]). The natural lengthscale is the OPM dispersal distance, so we set $\ell=0.5$km, given in [12] as the expected distance between a new nest and its parent nest. The wavespeed $v$ in Equation (2.2) can be interpreted as the expansion rate of the population and has units of km/year. Since $\ell=0.5$, we have $v=r$ km/year (with $r>0$, assumed implicitly in what follows). The specific forms we adopt for the growth rate $r$ are given in Section 2.3.
@@ 2.3     P roposed expansion rate(s)
for some real value $r>0$. The parametrisation of $M_{1}$ is $\theta_{1}=(r)$, where we adopt vector notation for consistency (see definition of $\theta_{2}$ below). This monophasic model $M_{1}$ encodes the assumption that the spread of OPM occurs at a constant rate $r$ throughout the period $2006$ to $2023$. The second model $M_{2}$ is of the form Equation (2.1) with
@@
where $r_{1},r_{2}>0$ are real values and the transition time $t_{c}$ satisfies $2006\leq t_{c}\leq 2023$. Hence, $M_{2}$ is parametrised by $\theta_{2}=(r_{1},r_{2},t_{c})$. This biphasic model $M_{2}$ encodes the assumption that the spread of OPM occurs at a non-constant rate; at a rate of $r_{1}$ from $2006$ up to and including $t_{c}$ and at a rate $r_{2}$ from $t_{c}$ to $2023$. We provide snapshots of a prototypical numerical solution to this model in Figure 3. We refer to these models as Model 1 and Model 2, respectively.
@@ 2.4     P arameter inference
The MCMC scheme (Gibbs sampler) described above is iterated until $m_{T}$ samples are obtained. We discard the initial $m_{B}\geq p$ samples as burn-in, and informally assess convergence of the remaining $m=m_{T}-m_{B}$ samples via visual diagnostics such as trace plots [26]. Additional scaling of the proposal variance to optimise acceptance rates can be carried out post hoc, if desired [41]. See Appendix A for more details.
@@ 2.5     M odel validation
We run the MCMC scheme for each of the models $M_{i}$ described above. To compare the models, we use the deviance information criterion (DIC), which measures relative quality of models [42]. For model $M_{i}$ (parametrised by $\theta_{i}$), the DIC is given by
@@
where $D(\theta_{i})=-2\log(\pi(\mathcal{D}|\theta_{i}))$ is the deviance of $\theta_{i}$; this criterion quantifies goodness of fit while penalising over-fitting. Models with a smaller DIC should be preferred to models with a larger DIC. We can readily compute $\bar{\theta}_{i}$ and $\overline{D(\theta_{i})}$ using the samples of $\theta$, obtained from the (final) MCMC run. Finally, we assess goodness of fit of the model with lowest DIC using posterior predictive sampling. Let $\tilde{t}_{k}$ denote a first arrival time at site $\mathbf{s}_{k}$. The predictive density is given by
@@ 3        R esults
Recall that Model 1 is parametrised by $\theta_{1}=(r)$ and $\sigma_{1}$. We initialise the MCMC scheme with values $\theta_{1,0}=4.0$ and $\sigma_{1,0}=2.0$, chosen arbitrarily. After burn-in and thinning the sample output to reduce auto-correlation, we obtain $m_{1}=10^{4}$ samples from the posterior distribution of the parameters of Model 1. This output is summarised by Figures 4 and 5, with the former suggesting satisfactory convergence of the MCMC scheme. In Table 2, we provide a summary of the inferred parameters for each model. For Model 1, with a constant expansion rate, we obtain a posterior mode of $\hat{r}=2.91$ with $95\%$ (equi-tailed) credible interval (CI) $(2.77,3.06)$ for the growth parameter $r$. For the observed noise parameter $\sigma$, the posterior mode is $\hat{\sigma}_{1}=4.85$ with $95\%$ CI $(4.54,5.21)$. The DIC reported for this model is $\text{DIC}_{1}=-1728.07$ (2dp).
@@ 4.1     T he spread of oak processionary moth in the UK
The values for the DICs of Models 1 and 2 (Table 2) indicate that Model 2 is the more appropriate choice for these observed data. This suggests that the expansion of oak processionary moth in the UK from 2006 to 2023 occurred at a non-constant rate. This conclusion aligns with previous results in the literature [9].
@@
Our evidence of biphasic expansion aligns with previous results, which also indicated two phases of expansion (at rates $1.66$ and $6.17$, respectively) with transition time $2014$ [9]. We note that the weighted average of $\hat{r}_{1}$ and $\hat{r}_{2}$ ($3.28$) is reasonably close in value to the Model 1 growth rate ($2.91$), and closer still to the mean of the previous biphasic estimates ($3.27$) [9]. Since the latter estimate covers $2006$ to $2019$, this could suggest that the expansion rate of OPM has continued to increase from $2019$ to $2023$. This is further supported by the increase in the phase 2 growth rate ($\hat{r}_{2}=6.21$ vs $6.17$). Nevertheless, we note that $6.17$ is within the 95CI of $r_{2}$, so it is possible the rate has stayed the same or decreased. In addition, methodological differences could explain these discrepancies.
@@ 4.2     M odelling approach
Previous estimates of the expansion rate of OPM in the UK were most recently obtained using a maximum distance method [9]. With our reaction-diffusion approach, we can obtain an estimate of the rate directly from our model parameters. This offers benefits; this framework can readily estimate future expansion rates. However, the equation for the expansion rate (Equation (2.2)) is obtained analytically, hence it will be susceptible to error introduced by the choice of numerical integration method. Fortunately, testing indicates acceptable convergence of the numerical scheme for our model, spatial-, and temporal-scales.
Our spatially-explicit modelling approach can provide an estimate of the arrival of OPM at every $1\text{km}$ area in a $256\text{km}^{2}$ spatial domain and can be readily extended to larger domain sizes, provided adequate computational resources are available. Our reaction-diffusion equation describes the evolution of the population density field; we apply a binary mask to this to facilitate comparison of our model with observed presence/absence data. Additionally, this could facilitate direct comparison of our approach with similar (compartmental) approaches [19, 20]. Furthermore, the population density field evolves with a clear outward-travelling wavefront (as in Figure 3). This captures the spatial trend of first arrival of the pest without incurring the computational cost of simulating a large number of individuals, for example. The use of a binary mask also allows our model to somewhat account for potential time lag in observation; time lag is posited to reflect the time needed for populations to reach detectable densities [46].
@@ 4.3     F urther Work
We thank Julia Branson from the University of Southampton and Andrew Hoppit from Forestry Commission England for useful discussions on the application to OPM, and the former for providing the OPM observational data. We also thank Kaitlyn N. Ries for helpful feedback on the manuscript.
   F unding
This work was completed as part of JPM’s PhD, supported by the Natural Environment Research Council (NERC) (Grant Number [NE/S007512/1]). LEW acknowledges support from a NERC Knowledge Exchange Fellowship (Grant Number [NE/X000478/1]).
   A uthor contributions
Jamie P. McKeown [conceptualization, methodology, software, validation, formal analysis, investigation, data curation, writing – original draft, writing – review and editing, visualization, project administration]; Laura E. Wadkin [conceptualization, methodology, data curation, writing – review and editing, supervision, project administration, funding acquisition]; Nick G. Parker [conceptualization, methodology, data curation, writing – review and editing, supervision, project administration, funding acquisition]; Andrew Golightly [methodology, writing – review and editing]; Andrew W. Baggaley [conceptualization, methodology, data curation, writing – review and editing, supervision, project administration, funding acquisition].
   D ata availability
The observational OPM nest data is available from Forestry Commission, England. Background maps of the UK used in Figures 1 and 2 are derived from oak abundance data in [27], available at [28]. Code and data produced by this work are available at [49].
@@ Appendix A Inferring model parameters
In step 3, the full conditional for $\theta$ is intractable, hence we adopt a random walk Metropolis-Hastings step to sample from the corresponding distribution. Define $\lambda=\log(\theta)$, where the logarithm is applied component-wise. In step 3 of the Gibbs sampler, we propose a new value $\lambda^{*}=\lambda+{\omega}$, where ${\omega}\sim N({0},\Sigma)$, for a strictly positive-definite innovation matrix $\Sigma$ – the covariance matrix of the proposal distribution. The choice of innovation matrix $\Sigma$ influences the mixing of the Markov chain. Practitioners often adopt an adaptive MCMC scheme wherein the innovation matrix $\Sigma$ is updated at fixed intervals to improve mixing [39]. We adopt the adaptive approach of Haario et al [39]. The MCMC scheme is instantiated with an arbitrary, strictly positive-definite, diagonal initial innovation matrix $\Sigma=\Sigma^{0}$ (we use the diagonal of the prior covariance matrix with zero covariances). We select a tuning period $p\ll m$ and a tuning interval $l\ll p$ where, for numerical convenience, we assume $p$ is divisible by $l$. Let $\theta^{(j)}$ denote the $j$th value in the chain. For the first $p$ samples, we update the innovation matrix at every $l^{\text{th}}$ sample. Hence the innovation matrix for the $j^{\text{th}}$ step of the MCMC scheme is given by
@@
where the scaling parameter $s_{d_{k}}=(2.38^{2})/{d_{k}}$ is a heuristic taken from [40] and $d_{k}$ is the dimension of the parameter space of model $M_{k}$. For a justification of this scheme, see [39].